OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

What's your birthday? Month and Day please.

OpenStudy (anonymous):

month: February Day 7th

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I figured out the answer to the first one @iGreen

OpenStudy (anonymous):

f(x)=3(x-2)^2+7

OpenStudy (igreen):

Lol you already have it, okay.

OpenStudy (igreen):

Standard form is: \(ax^2 + bx + c\)

OpenStudy (anonymous):

ok im gonna try to figure it out will post to make sure it's right i i think i have it

OpenStudy (igreen):

Okay.

OpenStudy (anonymous):

F(x)= 3x^2+12x-5

OpenStudy (igreen):

\(f(x)=3(x-2)^2+7\) First get rid of the exponent: \((x-2)^2 \rightarrow (x-2)(x-2)\) Multiply: \(x^2 - 2x - 2x + 4\) \(x^2-4x + 4\) Now plug that back into the equation: \(f(x)=3(x^2-4x+4)+7\) Now distribute the 3 into the parenthesis: \(f(x)=3x^2-12x+12+7\) Add: \(f(x)=3x^2-12x+19\)

OpenStudy (igreen):

You almost had it, the end is 19, not -5..I don't know how you made that mistake..

OpenStudy (igreen):

You probably subtracted 7 from 12 and made it negative.

OpenStudy (igreen):

So: Vertex Form: \(f(x)=3(x-2)^2+7\) is the same thing as: Standard Form: \(f(x)=3x^2−12x+19\)

OpenStudy (igreen):

If you see those black diamond question marks, just click on a different post and go back to this one and it'll go back to normal.

OpenStudy (anonymous):

OK thanks for the help on that, Green. now i dont understand the last one

OpenStudy (igreen):

Here's the graph:

OpenStudy (igreen):

Okay, first let's find x = 4 and x = 7: Plug x = 4 into the function and solve: \(f(x)=3(x-2)^2+7\) \(f(x)=3(4-2)^2+7\) Can you solve that? @Rocketman27

OpenStudy (igreen):

\(f(x)=3(4-2)^2+7\) Subtract: \(f(x)=3(2)^2+7\) Simplify exponent: \(f(x)=3(4)+7\) \(f(x)=12+7\) Add: \(f(x) = 19\) So we have (4, 19).

OpenStudy (igreen):

Now let's solve for x = 7: \(f(x)=3(x-2)^2+7\) \(f(x)=3(7-2)^2+7\) Subtract: \(f(x)=3(5)^2+7\) Simplify exponent: \(f(x)=3(25)+7\) Multiply: \(f(x)=75+7\) Add: \(f(x) = 82\) So we have (7, 82).

OpenStudy (igreen):

Now we have our two points, (4, 19) and (7, 82). To find the average rate of change between them, plug them into the slope equation, \(m = \dfrac{y_2-y_1}{x_2-x_1}\). In this case: \(y_2 = 82\) \(y_1 = 19\) \(x_2 = 7\) \(x_1 = 4\) Plug them in: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{82-19}{7-4}\) Subtract: \(m = \dfrac{63}{3}\) Divide: \(m = 21\) So the average rate of change between x = 4 and x = 7 is 21.

OpenStudy (igreen):

@Rocketman27

OpenStudy (anonymous):

Thank you. For the help @iGreen

OpenStudy (igreen):

No problem.